OpenStudy (anonymous):

A student opens a mathematics book to tow facing pages. The product of the page numbers is 1640. Find the page numbers. What is the number of the first page? What is the number of the second page?

OpenStudy (anonymous):

xy = 1640 x + 1 = y so plug: x(x+1) = 1640 ==> x^2 + x = 1640. quadratic equation to find x's (positive and negative value).